OpenStudy (anonymous):

Which fractions are equivalent to the fraction below? Check all that apply. 12/15 a. 4/5 b. 4/16 c. 24/30 d. 4/20

OpenStudy (anonymous):

12/15=0.8 check the other fractions ,which one equals 0.8

OpenStudy (texaschic101):

12/15 = 4/5 (a) 4/5 =4/5 (this one is equivalent) (b) 4/16 = 2/8 = 1/4 (not equivalent) (c) 24/30 = 4/5 (this one is equivalent) (d) 4/20 = 1/5 (not equivalent)